Can you solve 3(y- 5) is less than or equal to 6?
Answer by checkley77(12569)

3(y-5<=6
3y-15<=6
3y<=6+15
3y<=21
y<=21/3
y<=7 ans.
Proof:
Let y=6.9
3(6.9-5)<=6
3*1.9<=6
5.7<=6